OpenStudy (anonymous):

is anyone good at Adapting and Using Formulas?? PLEASE HELP question in the comments below??

OpenStudy (anonymous):

yes po .. @terenzreignz is good on that po, try to ask him about that po,, haha!

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

you're always welcome po. :D GOD bless. :D enjoy! haha!

OpenStudy (anonymous):

@terenzreignz please help?

OpenStudy (anonymous):

but theres no number to divide it by

OpenStudy (anonymous):

t=c(b+1) so divide by b+1

terenzreignz (terenzreignz):

Divide both sides of the equation by b+1, as permitted by the Multiplication Property of Equality.

terenzreignz (terenzreignz):

But wait, it's b you want to solve for...?

OpenStudy (anonymous):

im still not understanding sorry

OpenStudy (anonymous):

do you understand factoring out c as common factor

OpenStudy (anonymous):

no ive been on this test for atleast 2 days now its really hard

terenzreignz (terenzreignz):

Okay, I'm going to write the equation here... t = c + cb Right?

OpenStudy (anonymous):

so the question is looking for b \[t-c=bc\] \[b=\frac{t-c}{c}\]

OpenStudy (anonymous):